# Kandi, Sangareddy, Telengana 502285, India

### v1

@@ 1 Introduction
The AdS/CFT conjecture offers a non-perturbative framework for quantum gravity in asymptotically Anti-de Sitter (AdS) spacetime by connecting it to a boundary conformal field theory (CFT) Maldacena:1997re ; Witten:1998qj . It has become a central tool in advancing our understanding of quantum gravity, although a complete picture of holography is still lacking. A significant avenue of exploration lies in understanding how spacetime emerges from the underlying degrees of freedom of the field theory. In this context, the Ryu-Takayanagi (RT) proposal has played a particularly pivotal role in advancing our understanding Ryu:2006ef ; Hubeny:2007xt ; PhysRevD.34.373 ; Srednicki_1993 . The proposal states that the entanglement entropy (measure of quantum information) of a subregion $\mathbf{R}$ in the $d$-dimensional CFT is encoded in the area of a codimension-two minimal surface in the bulk, which is homologous to the boundary region $\mathbf{R}$.
@@
where $\cal{A}_{\mathbf{R}}$} is the area of the extremal surface anchored on the boundary of the subregion $\mathbf{R}$, and $G_{N}^{(d+1)}$ is the $(d+1)$ dimensional Newton’s constant. For time-dependent scenarios the proposal is generalized by Hubeny-Rangamani-Takayanagi (HRT) Hubeny:2007xt , which considers:
@@
While entanglement entropy captures essential features of spatial correlations, it is not suitable for describing the dynamics of the Einstein-Rosen (ER) bridge inside a black hole Susskind:2014moa . In holography, it is conjectured that the growth of the ER bridge in an AdS black hole corresponds to the increasing complexity of the boundary quantum state’s circuit Susskind:2014rva ; Stanford:2014jda . The computational complexity of a quantum state, defined as the number of unitary gates needed to reach the target state from an initial reference state, increases monotonically over time. This concept has led to the study of holographic complexity, considered dual to the quantum computational complexity of the boundary CFT state Susskind:2018pmk .
The holographic dual of circuit complexity in quantum computation theory has been shown to be linked to gravitational observables. These include the regularized volume of the largest codimension-one surface crossing the Einstein-Rosen bridges (“complexity = volume”(CV) proposal) Susskind:2014rva ; Stanford:2014jda , the low-energy on-shell action of the Wheeler-DeWitt (WDW) patch (“complexity = action” (CA) proposal) Brown:2015bva ; Brown:2015lvg , and the volume of the WDW patch (“ complexity = spacetime volume” (CV2.0) proposal) Couch:2016exn . Recently, it has been found that there is an infinite class of diffeomorphism-invariant gravitational observables that display universal features of complexity, which has led to the “complexity = anything” proposal Belin:2021bga ; Belin:2022xmt ; Jorstad:2023kmq . These proposals of holographic complexity have been studied in various holographic setups in AdS Carmi:2016wjl ; Carmi:2017jqz ; Chapman:2016hwi ; Alishahiha:2015rta ; Swingle:2017zcd ; Goto:2018iay ; An:2018dbz ; Jiang:2019qea ; Liu:2019mxz ; Jorstad:2023kmq ; Wang:2023ipy ; Zhang:2024mxb ; Emami:2024mes ; Omidi:2022whq ; Mandal:2022ztj ; Jiang:2025qai as well as (asymptotically) de Sitter (dS) spacetime Reynolds:2017lwq ; Chapman:2021eyy ; Aguilar-Gutierrez:2023pnn ; Aguilar-Gutierrez:2024rka ; Anegawa:2023wrk .
@@
In all the discussion above, it was assumed that the bulk gravitational theory is purely classical i.e. the central charge of the boundary CFT has been taken to be infinite. In this paper, we will take into account the quantum effects of bulk quantum fields on the holographic entanglement entropy as well as holographic complexity in the Karch-Randall (KR) braneworld setup Karch:2000ct , also known as double holography. Double holography facilitates us with three different descriptions of a theory.
  • Boundary Perspective: Describes a $d$-dimensional $\text{CFT}_{d}$ with a co-dimension one conformal defect [^1] The case when the $\text{CFT}_{d}$ has a boundary on which the (d1)(d-1)-dimensional CFT lives give rise to AdS/BCFT Takayanagi:2011zk .
@@
The double holography setup transforms the semiclassical problem of solving the field equations with quantum back-reactions to a classical problem by switching from the brane perspective to the bulk perspective. In other words, the classical black holes in $\text{AdS}_{d+1}$ with the brane describe quantum corrected black holes in $\text{AdS}_{d}$ with all orders of quantum corrections Emparan:2002px ; Emparan_2000 ; Emparan:1999fd .
In the context of entanglement entropy, the quantum corrections to the RT proposal have been found in Engelhardt_2015 ; Faulkner:2013ana
@@
where $A(\mathbf{X})$ is the area of the co-dimension two surface which is homologous to the boundary region $\mathbf{R}$ and $S_{\text{ent}}$ is the bulk entanglement entropy across the bulk extremal surface $\mathbf{X}$. It has been shown that this result can be understood using double holography Almheiri:2019hni , and the entanglement entropy is reproduced by considering the RT surface in the higher-dimensional bulk theory. In this setup of double holography, there are two possible RT surfaces connected (Hartman-Maldacena surface) and disconnected (island surfaces). The disconnected RT surface corresponds to the extremal surface that intersects with the brane and gives rise to the islands (a region on the brane inside the horizon). This led to the island proposal [^2] In KR setup due to the presence of bath, the AdS graviton become massive, for details on this refer toGeng:2020qvw Penington:2019npb ; Almheiri:2019psf , the entropy of some region $\mathbf{R}$ of a holographic quantum field theory that is coupled to a gravity theory is given by,
@@
While in Almheiri:2019hni they have considered 2-dimensional gravity+matter theory, the results have been generalized to eternal black holes in five-dimensional Schwarzschild AdS black holes Almheiri:2019psy , eternal charged Reissner-Nordström (RN) black holes in general dimensions Ling:2020laa , and dyonic black holes Jeong:2023lkc . Moreover, in Lin:2024gip , the entanglement entropy of a 4-dimensional Reissner-Nordström black hole has been studied, and the page curve is reproduced using the island formula. Also, a study of entanglement entropy where the bulk spacetime is pure $\text{AdS}_{d+1}$, and the boundary CFT lives on $\mathbb{R}\times S^{d-1}$ has been presented in Chen:2020uac .
In similar spirit, in the setup of braneworld holography, the quantum corrections to holographic complexity are explored for neutral black holes Emparan:2021hyr as well as rotating black holes Chen:2023tpi , following both CV and CA. The quantum corrections in the volume complexity can be distinguished into three different origins: (1) semiclassical backreaction on the classical black hole geometry $V(\Sigma)$, (2) higher derivative corrections to the gravity $v(\sigma)$, and (3) quantum corrections due to bulk quantum fields $C_{V}^{\text{bulk}}(\lvert\psi\rangle)$, while such a distinction cannot be achieved in the case of action complexity. Still, the action conjecture has been much more efficient in the braneworld setup, as it provides holographic complexity incorporating all order corrections due to bulk quantum fields.
@@
In this work, we build upon previous investigations of holographic complexity in neutral and rotating black holes Emparan:2021hyr ; Chen:2023tpi . We extend the analysis to dyonic charged quantum black hole geometries, introducing an unexplored probe of holographic entanglement entropy within the setup that has not been previously investigated. Another motivation to consider this system is that it serves as a practical model for exploring magneto-transport phenomena and quantum Hall states in strongly coupled quantum theories that operate under different chemical potentials and curved background geometries. These aspects have wide applications in condensed matter physics. More broadly, this framework allows for the study of defects that can carry both electric charge and magnetic flux, which could potentially model anyonic excitations.
@@ 2 Quantum charged black holes
In this section, we present the main features of the quantum charged black hole. A three-dimensional quantum charged black hole can be constructed through the KR braneworld approach by placing an $\text{AdS}_{3}$ brane with tension $\tau$ within the four-dimensional bulk spacetime described by a four-dimensional AdS charged C-metric. Following the notation of Climent:2024nuj ; Feng:2024uia , the total action of the system is expressed as:
@@ 2.1 Review of the charged AdS C-metric & thermal quantities
In this section, we outline the key characteristics of the charged extension of the quantum BTZ (quBTZ) black hole Climent:2024nuj ; Feng:2024uia . The $\mathrm{AdS}_{4}$ C-metric, which includes both electric and magnetic charges of the bulk black hole, originates from a four-dimensional solution of the Einstein-Maxwell-AdS theory. This four-dimensional setup describes a three-dimensional quantum black hole on a brane placed in the bulk within the context of braneworld holography. The metric of the four-dimensional charged AdS black hole can be conveniently written in the following form:
@@
We place a brane at $x=0$ in the bulk spacetime (10), and truncate the bulk at the position of the brane. Then, to complete the spacetime, we glue an identical copy of the bulk at the brane position. The Israel junction condition across the brane $K^{+}_{\mu\nu}-K^{-}_{\mu\nu}=\frac{\partial S}{\partial h_{\mu\nu}}$ fix the brane tension Israel:1966rt
@@
This metric can be seen as a solution to the three dimensional effective theory of gravity coupled to the gauge field on the brane Climent:2024nuj , and therefore it is interpreted as the quantum backreacted black holes, often referred as “quantum black holes”.
@@
where $\mathcal{G}_{3}$ is ‘renormalised’ three dimensional Newton’s constant, given by
@@
and $M$ is the mass of the black hole. Following Kudoh:2004ub ; Emparan_2006 ; Climent:2024nuj ; Feng:2024uia , one can express the mass in terms of the conical deficit as [^6] Notice that for the case of black holes i.e. $\kappa=-1$, $M>0$ and for the quantum dressed conical defects i.e. $\kappa=1$ the mass is negative $M<0$
@@
In addition to the standard thermodynamic quantities, it has been shown in Kubiznak:2012wp ; Kubiznak:2016qmn ; Frassino:2022zaz that one can also introduce a notion of thermodynamic pressure, defined as
@@
Moreover, if the brane tension is treated as a thermodynamic variable by allowing variations in the backreaction parameter $\ell$, one obtains a corresponding thermodynamic area of the brane, given by
@@ 3 Entanglement Entropy
To study the disconnected RT surface, we again consider the same boundary region of radius $\epsilon$ around $x_{1}$ as we did for the connected RT surface. While for the disconnected RT suface we have divided the RT surface between two regions from $\{0,x_{c}\}$ and $\{r_{c},r_{b}\}$, with $x(r_{c})=x_{c}$ and vice versa, following Almheiri:2019psy . This uplifts the ambiguity if the RT surface is multi-valued, and the area functional is given by,
@@
Figure 4: (a) $A_{D}$ vs $r_{0}$ for $\kappa=-1$ and $q=0$ (b) A zoomed in version of $A_{D}$ vs $r_{0}$ for $\kappa=-1$ and $q=0$, (c) A zoomed in version of $A_{D}$ vs $r_{0}$ for $\kappa=-1$ and different non-zero $q$’s
@@
To fix the parameter $\alpha$, we study the behavior of the area functional with $\alpha$ as shown in figure 3, and choose the value of $\alpha$ which minimizes the area functional. This leaves both area functional $A_{D}$ and $r_{0}$ in terms of $x_{c}$ and $r_{c}$ at the fixed value of $\alpha$, so that we can study the behavior of the entanglement entropy of the disconnected phase as a function of $r_{0}$, this has been shown in figure 4 for $\kappa=-1$ and in figure 5 for $\kappa=1$.
Figure 5: (a) $A_{D}$ vs $r_{0}$ for $\kappa=1$ and $q=0$ (b) A zoomed out version of $A_{D}$ vs $r_{0}$ for $\kappa=1$ and $q=0$, (c) A zoomed out version of $A_{D}$ vs $r_{0}$ for $\kappa=1$ and different non-zero $q$’s
As shown in figure 4 and 5, the area functional attains a minimum with respect to $r_{0}$ for both $\kappa=\pm 1$. Therefore as $r_{0}$ determines the intersection point with the brane (or the island region), this depicts the validity of the island proposal for the quantum black holes.

### v2

# Kandi, Sangareddy, Telangana 502285, India

@@ 1 Introduction
The AdS/CFT conjecture offers a non-perturbative framework for quantum gravity in asymptotically Anti-de Sitter (AdS) spacetime by connecting it to a boundary conformal field theory (CFT) Maldacena (1998); Witten (1998). It has become a central tool in advancing our understanding of quantum gravity, although a complete picture of holography is still lacking. A significant avenue of exploration lies in understanding how spacetime emerges from the underlying degrees of freedom of the field theory. In this context, the Ryu-Takayanagi (RT) proposal has played a particularly pivotal role in advancing our understanding Ryu and Takayanagi (2006); Hubeny et al. (2007); Bombelli et al. (1986); Srednicki (1993). The proposal states that the entanglement entropy (measure of quantum information) of a subregion $\mathbf{R}$ in the $d$-dimensional CFT is encoded in the area of a codimension-two minimal surface in the bulk, which is homologous to the boundary region $\mathbf{R}$.
@@
where $\cal{A}_{\mathbf{R}}$} is the area of the extremal surface anchored on the boundary of the subregion $\mathbf{R}$, and $G_{N}^{(d+1)}$ is the $(d+1)$ dimensional Newton’s constant. For time-dependent scenarios the proposal is generalized by Hubeny-Rangamani-Takayanagi (HRT) Hubeny et al. (2007), which considers:
@@
While entanglement entropy captures essential features of spatial correlations, it is not suitable for describing the dynamics of the Einstein-Rosen (ER) bridge inside a black hole Susskind (2016b). In holography, it is conjectured that the growth of the ER bridge in an AdS black hole corresponds to the increasing complexity of the boundary quantum state’s circuit Susskind (2016a); Stanford and Susskind (2014). The computational complexity of a quantum state, defined as the number of unitary gates needed to reach the target state from an initial reference state, increases monotonically over time. This concept has led to the study of holographic complexity, considered dual to the quantum computational complexity of the boundary CFT state Susskind (2018).
The holographic dual of circuit complexity in quantum computation theory has been shown to be linked to gravitational observables. These include the regularized volume of the largest codimension-one surface crossing the Einstein-Rosen bridges (“complexity = volume”(CV) proposal) Susskind (2016a); Stanford and Susskind (2014), the low-energy on-shell action of the Wheeler-DeWitt (WDW) patch (“complexity = action” (CA) proposal) Brown et al. (2016b, a), and the volume of the WDW patch (“ complexity = spacetime volume” (CV2.0) proposal) Couch et al. (2017). Recently, it has been found that there is an infinite class of diffeomorphism-invariant gravitational observables that display universal features of complexity, which has led to the “complexity = anything” proposal Belin et al. (2022, 2023); Jørstad et al. (2023). These proposals of holographic complexity have been studied in various holographic setups in AdS Carmi et al. (2017b, a); Chapman et al. (2017); Alishahiha (2015); Swingle and Wang (2018); Goto et al. (2019); An et al. (2018); Jiang and Zhang (2020); Liu et al. (2020); Jørstad et al. (2023); Wang and Ren (2023); Zhang et al. (2024); Emami and Parvizi (2025); Omidi (2023); Mandal et al. (2022); Jiang and Liu (2025); Jiang et al. (2024); Wang et al. (2023) as well as (asymptotically) de Sitter (dS) spacetime Reynolds and Ross (2017a); Chapman et al. (2022); Aguilar-Gutierrez (2024); Aguilar-Gutierrez et al. (2024a); Anegawa et al. (2023); Paul et al. (2025); Susskind (2021); Jørstad et al. (2022); Baiguera et al. (2023); Anegawa and Iizuka (2023); Aguilar-Gutierrez et al. (2024c).
@@
In all the discussion above, it was assumed that the bulk gravitational theory is purely classical i.e. the central charge of the boundary CFT has been taken to be infinite. In this paper, we will take into account the quantum effects of bulk quantum fields on the holographic entanglement entropy [^1] A similar study has been done in Chakrabortty et al. (2020); Pant et al. (2025) where the authors have consider the quantum backreaction by heavy-flavor insertions and boundary deformations on the HEE, HC and choas. as well as holographic complexity in the Karch-Randall (KR) braneworld setup Karch and Randall (2001), also known as double holography. Double holography [^2] A top-down string theory realization of double holography has been given in Uhlemann (2021); Karch et al. (2022); Deddo et al. (2024). facilitates us with three different descriptions of a theory.
  • Boundary Perspective: Describes a $d$-dimensional $\text{CFT}_{d}$ with a co-dimension one conformal defect [^3] The case when the $\text{CFT}_{d}$ has a boundary on which the (d1)(d-1)-dimensional CFT lives give rise to AdS/BCFT Takayanagi (2011)..
@@
The double holography setup transforms the semiclassical problem of solving the field equations with quantum back-reactions to a classical problem by switching from the brane perspective to the bulk perspective. In other words, the classical black holes in $\text{AdS}_{d+1}$ with the brane describe quantum corrected black holes in $\text{AdS}_{d}$ with all orders of quantum corrections Emparan et al. (2002, 2000a, 2000b).
In the context of entanglement entropy, the quantum corrections to the RT proposal have been found in Engelhardt and Wall (2015); Faulkner et al. (2013)
@@
where $A(\mathbf{X})$ is the area of the co-dimension two surface which is homologous to the boundary region $\mathbf{R}$ and $S_{\text{ent}}$ is the bulk entanglement entropy across the bulk extremal surface $\mathbf{X}$. It has been shown that this result can be understood using double holography Almheiri et al. (2020a), and the entanglement entropy is reproduced by considering the RT surface in the higher-dimensional bulk theory. In this setup of double holography, there are two possible RT surfaces connected (Hartman-Maldacena surface) and disconnected (island surfaces). The disconnected RT surface corresponds to the extremal surface that intersects with the brane and gives rise to the islands (a region on the brane inside the horizon). This led to the island proposal [^4] In KR setup due to the presence of bath, the AdS graviton become massive, for details on this refer to Geng and Karch (2020). Penington (2020); Almheiri et al. (2019), the entropy of some region $\mathbf{R}$ of a holographic quantum field theory that is coupled to a gravity theory is given by,
@@
While in Almheiri et al. (2020a) they have considered 2-dimensional gravity+matter theory, the results have been generalized to eternal black holes in five-dimensional Schwarzschild AdS black holes Almheiri et al. (2020b), eternal charged Reissner-Nordström (RN) black holes in general dimensions Ling et al. (2021), and dyonic black holes Jeong et al. (2023). Moreover, in Lin et al. (2024), the entanglement entropy of a 4-dimensional Reissner-Nordström black hole has been studied, and the page curve is reproduced using the island formula. Also, a study of entanglement entropy where the bulk spacetime is pure $\text{AdS}_{d+1}$, and the boundary CFT lives on $\mathbb{R}\times S^{d-1}$ has been presented in Chen et al. (2020).
In a similar spirit, in the setup of braneworld holography, the quantum corrections to holographic complexity [^5] Also in Hernandez et al. (2021) the authors have studied subregion complexity in a double holographic setup. are explored for neutral black holes Emparan et al. (2022) as well as rotating black holes Chen et al. (2024), following both CV and CA. The quantum corrections in the volume complexity can be distinguished into three different origins: (1) semiclassical backreaction on the classical black hole geometry $V(\Sigma)$, (2) higher derivative corrections to the gravity $v(\sigma)$, and (3) quantum corrections due to bulk quantum fields $C_{V}^{\text{bulk}}(\lvert\psi\rangle)$, while such a distinction cannot be achieved in the case of action complexity. Still, the action conjecture has been much more efficient in the braneworld setup, as it provides holographic complexity incorporating all order corrections due to bulk quantum fields.
@@
In this work, we build upon previous investigations of holographic complexity in neutral and rotating black holes Emparan et al. (2022); Chen et al. (2024). We extend the analysis to dyonic charged quantum black hole geometries, introducing an unexplored probe of holographic entanglement entropy within the setup that has not been previously investigated. Another motivation to consider this system is that it serves as a practical model for exploring magneto-transport phenomena and quantum Hall states in strongly coupled quantum theories that operate under different chemical potentials and curved background geometries. These aspects have wide applications in condensed matter physics. More broadly, this framework allows for the study of defects that can carry both electric charge and magnetic flux, which could potentially model anyonic excitations.
@@ 2 Quantum charged black holes
In this section, we present the main features of the quantum charged black hole. A three-dimensional quantum charged black hole can be constructed through the KR braneworld approach by placing an $\text{AdS}_{3}$ brane with tension $\tau$ within the four-dimensional bulk spacetime described by a four-dimensional AdS charged C-metric. Following the notation of Climent et al. (2024); Feng et al. (2024), the total action of the system is expressed as:
@@ 2.1 Review of the charged AdS C-metric & thermal quantities
In this section, we outline the key characteristics of the charged extension of the quantum BTZ (quBTZ) black hole Climent et al. (2024); Feng et al. (2024). The $\mathrm{AdS}_{4}$ C-metric, which includes both electric and magnetic charges of the bulk black hole, originates from a four-dimensional solution of the Einstein-Maxwell-AdS theory. This four-dimensional setup describes a three-dimensional quantum black hole on a brane placed in the bulk within the context of braneworld holography. The metric of the four-dimensional charged AdS black hole can be conveniently written in the following form:
@@
We place a brane at $x=0$ in the bulk spacetime (10), and truncate the bulk at the position of the brane. Then, to complete the spacetime, we glue an identical copy of the bulk at the brane position. The Israel junction condition across the brane $K^{+}_{\mu\nu}-K^{-}_{\mu\nu}=\frac{\partial S}{\partial h_{\mu\nu}}$ fix the brane tension Israel (1966)
@@
This metric can be seen as a solution to the three dimensional effective theory of gravity coupled to the gauge field on the brane Climent et al. (2024), and therefore it is interpreted as the quantum backreacted black holes, often referred as “quantum black holes”.
@@
where $\mathcal{G}_{3}$ is ‘renormalized’ three-dimensional Newton’s constant, given by
@@
and $M$ is the mass of the black hole. Following Kudoh and Kurita (2004); Emparan (2006); Climent et al. (2024); Feng et al. (2024), one can express the mass in terms of the conical deficit as [^9] Notice that for the case of black holes i.e. $\kappa=-1$, $M>0$ and for the quantum dressed conical defects i.e. $\kappa=1$ the mass is negative $M<0$.
@@
In addition to the standard thermodynamic quantities, it has been shown in Kubiznak and Mann (2012); Kubiznak et al. (2017); Frassino et al. (2023) that one can also introduce a notion of thermodynamic pressure, defined as
@@
Moreover, if the brane tension is treated as a thermodynamic variable by allowing variations in the backreaction parameter $\ell$, one obtains a corresponding thermodynamic area of the brane given by,
@@ 3 Entanglement Entropy
To study the disconnected RT surface, we again consider the same boundary region of radius $\epsilon$ around $x_{1}$ as we did for the connected RT surface. While for the disconnected RT suface we have divided the RT surface between two regions from $\{0,x_{c}\}$ and $\{r_{c},r_{b}\}$, with $x(r_{c})=x_{c}$ and vice versa, following Almheiri et al. (2020b). This uplifts the ambiguity if the RT surface is multi-valued, and the area functional is given by,
@@
Figure 4: (a) The area profile of the disconnected surface $A_{D}$ is shown as a function of the intersection point $r_{0}$, with $x_{c}$ varied and $r_{c}$ and $\alpha$ minimized, for $\kappa=-1$ and $q=0$. (b) The area profile of $A_{D}$ is plotted as a function of $r_{0}$, varying $\alpha$ while minimizing $x_{c}$ and $r_{c}$, exhibits a shallow dip. (c) The plots in (b) are now shown for several non-zero values of charges.
@@
Among these three parameters, we first fix the parameter $\alpha$, so that the area is minimized. The explicit behavior of the area with $\alpha$ is shown in figure 3. This leaves us with $A_{D}$ as a function of two remaining parameters. Furthermore we observe that the area is a monotonic function of the parameters $x_{c}$ and $r_{c}$, this enables us to fix either one of them so that area is minimized and we can study the parametric dependence of area $A_{D}$ on the intersection point $r_{0}$ by varying the other parameter. In figure 4 and 5 we have shown a parametric plot of $A_{D}(r_{0})$ by varying $x_{c}$ and fixing $(\alpha,r_{c})$ at their minimizing values. For both signs of $\kappa$, the resulting curve is almost monotonic on the full range of $r_{0}$, and we do not see any explicit minima in these plots. But, if we choose to fix the parameters $(x_{c},r_{c})$ and study the parametric plot with respect to $\alpha$, this enables us to see a minima explicitly, they are presented [^13] The plots shown in figure 4-5(a) and 4-5(b) are different parametric slices of $A_{D}(x_{c},r_{c},\alpha)$ and need not to match pointwise just they must agree near the extremum point given in terms of $r_{0}(x_{c},r_{c},\alpha)$. in figure 4 and 5.
Figure 5: (a) The area profile of the disconnected surface $A_{D}$ is shown as a function of $r_{0}$, with $x_{c}$ varied and $r_{c}$ and $\alpha$ minimized, for $\kappa=1$ and $q=0$. (b) The area profile of $A_{D}$ is plotted as a function of $r_{0}$, varying $\alpha$ while minimizing $x_{c}$ and $r_{c}$. (c) The plots in (b) shown for several non-zero values of $q$.
At last, in figures 4 and 5, we have shown the parametric behavior of $A_{D}$ with the intersection point $r_{0}$ for different values of charge parameters $q$, emphasizing how charge shifts the minima. As shown in figure 4 and 5, the area functional attains a minimum with respect to $r_{0}$ for both $\kappa=\pm 1$. Therefore as $r_{0}$ determines the intersection point with the brane (or the island region), this depicts the validity of the island proposal for the quantum black holes.